OpenStudy (anonymous):

use the elimination method to solve the system of equations 5u+2v=-18 3u+v=-6

OpenStudy (anonymous):

Would you like to get rid of the u or the v? One is easier than the other.

OpenStudy (anonymous):

@jabberwock i just tying to solve the problem

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

Alright, let's get rid of the v. You have 2v in the first equation and v in the second. You want the coefficients to be opposites of each other. So you want to multiply everything in the second equation by -2.

OpenStudy (uri):

5u+2v=-18 3u+v=-6 Multiply the 2nd eq by -2 So, 5u+2v=-18 -6u-2y=12 ========= Solve this.

OpenStudy (anonymous):

2y? @uri

OpenStudy (uri):

It's 2v*

OpenStudy (anonymous):

1u=6

OpenStudy (uri):

Always remember the sign of greater number comes so it's -6+5=-1 and -18+12=-6 So,-1u=-6

OpenStudy (uri):

Now divide -6 with -1

OpenStudy (anonymous):

-6

OpenStudy (uri):

Aha..-6/-1 double negative=positive. So,u=6

OpenStudy (anonymous):

@uri right u=6

OpenStudy (uri):

Now we have u,Lets solve for v,Put the value of u i.e 6 in original eq 2..So, 3u+v=-6 3(6)+v=-6 Solve it.

OpenStudy (anonymous):

18+v=-6, v=3

OpenStudy (anonymous):

@uri

OpenStudy (uri):

3(6)+v=-6 18+v=-6 v=-6-18 v=-24.